# Context and Diversity Matter: The Emergence of In-Context Learning in World Models

### v1

@@ 3.4 L2World: Long-Context and Linear-Attention World Models
Prior work relies on multi-token representations or diffusion models to deliver high-fidelity single-frame reconstructions specifically for images. While these approaches set the state-of-the-art for static images, they introduce prohibitive memory and computational bottlenecks when sequences grow to the length required for ICEL. We therefore introduce L2World that trade per-frame fidelity for temporal scalability: for image observations, we compress each frame $o_{t}$ into a latent state $s_{t}$ with a lightweight variational auto-encoder (VAE) [66] whose encoder $f_{I}$ and decoder $g_{I}$ are ResNet stacks; for low-dimensional observations, we simply apply a small multi-layer perceptron encoder/decoder pair. For computational efficiency, we do not model the full state estimation with $\hat{p}(s_{t}|q_{t})$. Instead, we construct a pseudo-state that depends solely on the instant observation and leaves all temporal-related encoding to transition modeling [4]. We then construct the adaptive world model $\hat{p}_{\theta}(o_{t+1}|q_{t},C_{T})$ using an efficient sequence decoder $f_{\theta}$. Here we implement gated slot attention layers [67, 68, 69] with chunk-wise parallelization during the training phase, while retaining the recurrent form during the inference phase. The predictor of transition at first yields the output $h_{t}$, which is further processed by the decoder $g_{S}$ to produce the predicted state $\hat{s}$, corresponding to predicting a Gaussian distribution over the latent space $\hat{p}\sim\mathcal{N}(\hat{s}_{t},\sigma^{2}_{s})$. Although this assumption could lead to significant loss of accuracy in stochastic environments, for the navigation tasks we consider, it is acceptable and greatly increases computational efficiency. The model and the target function are listed as follows (see details in Section˜B.2):
@@ 4 Experiments
Although prior studies [62, 65, 53] have validated the influence of data distribution on ICL, they have largely concentrated on simplified tasks such as regression and classification. To examine how data distribution, model architecture, and training procedure jointly affect the emergence of ICEL/ICER, we select two canonical benchmarks. First, the cart-pole, a classical continuous-control problem, in which ICEL primarily targets the acquisition of varying embodiments and physical constants. Second, indoor navigation, a widely recognized POMDP, in which ICEL focuses on learning and memorizing spatial coefficients. These two experiments confirm not only the impact of each factor on ICEL but also demonstrate that ICEL spans a broad generalization spectrum, extending from spatial reasoning and memorization to adaptation of embodiment and physical constants.
@@ 4.1 Random Cart-Poles
Experiment Setting. To investigate ICEL and ICER in cart-pole environments, we randomized four variables in the environment settings: gravity $g$, cart mass $m_{c}$, pole mass $m_{p}$, and pole length $l$. We focus on two different scopes of the configurations to investigate the impact of the diversity issue: Scope 1 remains close to the original task, whereas Scope 2 covers a larger region and excludes Scope 1 (details are left to Figure˜5). For each environment, we first trained an RL agent and then collected trajectories with an expert policy perturbed by uniform noise spanning $[0.3,0.7]$ to ensure adequate coverage of the state-action space. We trained five comparison models; all share the same data scale (128K trajectories $\times$ 200 step/trajectory) but differ in the number and scope of the environments.
@@
Figure 2: Comparison of models trained on different datasets (color-coded) in Cart-Poles. Performance varies markedly with the training data, revealing distinct tendencies toward ICER, ICEL, or an inability to perform ICL.
@@
Importance of both environment scope and environment number: A comparison between the model trained on 1 environment (1 Env) and 4 environments (4 Envs) with the other group demonstrates that an insufficient number of environments leads to the absence of ICL and generalization, except in the tasks that the model has already seen. The 4 Envs group exhibits clear ICER characteristics, with a substantial performance gap between seen and unseen tasks. The 16 Envs (Scope 1 + 2) and 8K Envs (Scope 1) groups display similar capabilities; however, they lag significantly behind the 8K Envs (Scope 1 + 2) group, indicating that both the scope of tasks and the number of tasks are crucial.
@@ 4.2 Navigation
Experiment settings. Procedurally generated mazes are a demanding test-bed for transition prediction [70, 71]. By stripping observations of semantic cues, the maze framework keeps task complexity low while still exposing models to stochastic, partially observable dynamics. We amplify diversity through fully randomized configurations that vary topology, textures, object placement, and agent embodiment  [^1] Procedural Maze environments:
@@
. Room-tour data are collected in a way similar to cart-pole: we perturb the oracle object-navigation policy [72, 70] with random noise levels in $[0.05,0.95]$ and record the complete trajectory of an agent in each environment. The oracle policy was derived using Dijkstra’s algorithm based on the ground truth 2-D occupancy maps. Observations are RGB images standardized to 128 × 128 pixels; The action space comprises 17 discrete actions, each corresponding to a unique offset and rotational movement. Table 1 lists the resulting training sets, each drawn from a different number of environments and exhibiting varying trajectory lengths. To isolate the effect of data distribution, all maze datasets contain the same number of frames while differing in the coverage of trajectories and environments. To further investigate its transferability to more realistic environments, we also collect trajectories from the semantically rich ProcTHOR simulation, which offers a wide variety of assets [73, 74]. Specifically, we curate two datasets: a larger one with 40K trajectories and a smaller one with 5K trajectories, each trajectory having a length of 2K.
@@
For training, inspired by overshooting [43] and Hu et al. [38], we randomly mask the $s_{t}$ of the input sequences at some positions to enhance the model’s capability to predict the distant future (see training details in Section˜B.1). Evaluation is conducted on both seen and unseen tasks. By default, we use an evaluation set scale of $|\mathcal{E}|=256$. The evaluation process involves encoding a context of length $t$ for ICEL and then predicting future $k$-step transitions using auto-regression and the ground truth action records $a_{t+1},a_{t+2},...,a_{t+k}$. We assess the error in both the latent spaces and the decoded images, which we refer to as $k$-step prediction with context length $T=t$. For example, $T=10$ and $k=4$ predict future $4$ steps with a context length of $10$.
We evaluate two additional baselines alongside the proposed long-context world model: (1) Navigation World Model (NWM) [7], which employs diffusion layers to predict the next frame from the preceding four frames; (2) Dreamer-v3 [34], which uses LSTM layers for temporal encoding. For NWM, we retain its original pre-trained image encoders and re-train only the diffusion layers on the target dataset. For Dreamer-v3, we remove the policy components and train only the world-model module to ensure a fair comparison.
@@
Impact of data distribution and model architecture on ICEL. Table˜2 reports the next-frame prediction quality estimated by PSNR ($k=1$) for models trained on different datasets. Three empirical findings corroborate Theorem˜1: (1) The 32K-L dataset yields the best generalization to unseen environments, whereas the 128-L dataset excels on seen ones; in both cases, peak performance occurs at the asymptotic stage, not at the beginning of the context. (2) Long-context training consistently produces stronger ICL than short-context training, confirming that extensive context is necessary for ICL to emerge. (3) Dreamer and NWM fall short even with a long-context dataset: Dreamer’s LSTM backbone and NWM’s 4-frame horizon show that architectures incapable of fully leveraging long contexts cannot achieve many-shot ICEL. Figure˜3 further presents the $k=\{1,2,4,8\}$-step prediction performances on unseen Mazes, measuring how far ahead the world models can reliably foresee. For $k>1$, the performance–context-length curve largely tracks the next-frame trend across models, except for Dreamer (Maze 32K-L): at $k=8$, its performance plateaus once $T>100$, whereas $k=1$ keeps improving, revealing a larger compound-error accumulation than in our method.
Figure˜4 illustrates 10-step-ahead predictions at $T=\{1,100,10K\}$ for our method, Dreamer, and NWM trained on Maze-32K-L. NWM produces visually convincing frames with fine textures and a plausible layout, yet frame-wise fidelity alone is insufficient for correct long-range forecasting because the model lacks long-term memory and spatial reasoning.
@@
Transferability of ICEL and ICER: To verify whether ICEL emerges in semantically richer, realistic environments and whether skills acquired in simpler environments transfer to more complex ones, we fine-tune the models on ProcTHOR trajectories and evaluate them in varied contexts, as shown in Table˜3. The model pre-trained on Maze-32K-L not only excels in unseen mazes but also retains its advantage when fine-tuned on the small ProcTHOR-5K set; its transferability markedly exceeds that of Maze-128-L and other baselines, evidencing ICEL’s domain-generality. Further increasing ProcTHOR data yields continual improvement in our model while maintaining a large margin over Dreamer and NWM. However, the performance from $T=1K$ to $T=10K$ begins to deteriorate once the training data switches from ProcTHOR-5K to ProcTHOR-40K, implying that ProcTHOR trajectories with $T\leq 2K$ fail to incentivize ICEL at longer contexts. Yet, this phenomenon does not appear with ProcTHOR-5K, verifying that the ICEL capability acquired in L2World transfers effectively to ProcTHOR despite the large gap between the two classes of environments.
@@ 5 Conclusions and Limitations
Conclusions: This work investigates in-context learning of world models, specifically dynamic models, focusing on the possible modes of ICEL and ICER in MDP and POMDP. Theoretically, we analyze error upper bounds for both modes to characterize their properties and identify the conditions under which each excels. Empirically, we introduce L2World and validate these insights in cart-pole and navigation tasks. Our results underscore that both high environment diversity and sufficient context length of the world model are essential to elicit ICEL.
@@
Acknowledgments and Disclosure of Funding
This project is supported by Longgang District Shenzhen’s “Ten Action Plan” for supporting innovation projects (under Grant LGKCSDPT2024002).
@@ Appendix A Assumptions and Proofs for Theorem˜1
Proof for the first part of Theorem˜1: We first define
@@
First, we gave that
@@
$=TV(\hat{p}_{\theta,\hat{e}},\hat{p}_{\theta,\hat{e}_{0}})+\min_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}})$
@@
$<\frac{\alpha(|\mathcal{E}|-1)}{3\sqrt{T}}$
@@
To estimate $\hat{p}_{EL}(s^{\prime}|s,a)$, we use Equation˜4 to acquire:
@@
Now employ Equation˜6 and Equation˜7 in the calculation of TV:
@@ B.1 Environments and Datasets
Cart-pole: Figure˜5 lists the scopes of the Cart-Pole environment variants, their t-SNE visualization, and details of the training and evaluation data. All datasets share a trajectory length of 200, which is sufficient for ICL in Cart-Pole variants. The training data comprises a total of 25.6M timesteps to avoid interference from data scale on performance, and the evaluation data contains 205K steps in total.
@@
Mazes: Our maze environment is closely related to the settings described in Pašukonis et al. [70], Wang et al. [75], which are generated on a $15\times 15$ grid world. The distribution of the sampled mazes is shown in Figure˜6 and Figure˜7. The primary distinction between our mazes and those in previous work is the enhanced environmental diversity achieved through randomized configurations, which include the following:
@@
Figure 8: Illustration of the embodiment, the observation and the trajectories in procedurally generated mazes (a) and ProcTHOR (b).

### v2

@@ 3.4 L2World: Long-Context and Linear-Attention World Models
Prior work relies on multi-token representations or diffusion models to deliver high-fidelity single-frame reconstructions specifically for images. While these approaches set the state-of-the-art for static images, they introduce prohibitive memory and computational bottlenecks when sequences grow to the length required for EL. We therefore introduce L2World that trades per-frame fidelity for temporal scalability: for image observations, we compress each frame $o_{t}$ into a latent state $s_{t}$ with a lightweight variational auto-encoder (VAE) (Kingma and Welling, 2014) whose encoder $f_{I}$ and decoder $g_{I}$ are ResNet stacks; for low-dimensional observations, we simply apply a small multi-layer perceptron encoder/decoder pair. For computational efficiency, we do not model the full state estimation with $\hat{p}(s_{t}|q_{t})$. Instead, we construct a pseudo-state that depends solely on the instant observation and leaves all temporal-related encoding to transition modeling Mazzaglia et al. (2024). We then construct the adaptive world model $\hat{p}_{\theta}(o_{t+1}|q_{t},C_{T})$ using an efficient sequence decoder $f_{\theta}$. Here we implement gated slot attention layers Wang et al. (2020); Yang et al. (2024); Zhang et al. (2024b) with chunk-wise parallelization during the training phase, while retaining the recurrent form during the inference phase. The predictor of transition at first yields the output $h_{t}$, which is further processed by the decoder $g_{S}$ to produce the predicted state $\hat{s}$, corresponding to predicting a Gaussian distribution over the latent space $\hat{p}\sim\mathcal{N}(\hat{s}_{t},\sigma^{2}_{s})$. Although this assumption could lead to significant loss of accuracy in stochastic environments, for the navigation tasks we consider, it is acceptable and greatly increases computational efficiency. The model and the target function are listed as follows (see details in Section˜B.2):
@@ 4 Experiments
Although prior studies Chan et al. (2022); Singh et al. (2023); Raventós et al. (2023) have validated the influence of data distribution on ICL, they have largely concentrated on simplified tasks such as regression and classification. To examine how data distribution, model architecture, and training procedure jointly affect the emergence of EL/ER, we select two canonical benchmarks. First, the cart-pole, a classical continuous-control problem, in which EL primarily targets the acquisition of varying embodiments and physical constants. Second, indoor navigation, a widely recognized POMDP, in which EL focuses on learning and memorizing spatial coefficients. These two experiments confirm not only the impact of each factor on EL but also demonstrate that EL spans a broad generalization spectrum, extending from spatial reasoning and memorization to adaptation of embodiment and physical constants.
@@ 4.1 Random Cart-Poles
Experiment Setting. To investigate EL and ER in cart-pole environments, we randomized four variables in the environment settings: gravity $g$, cart mass $m_{c}$, pole mass $m_{p}$, and pole length $l$. We focus on two different scopes of the configurations to investigate the impact of the diversity issue: Scope 1 remains close to the original task, whereas Scope 2 covers a larger region and excludes Scope 1 (details are left to Figure˜6). For each environment, we first trained an RL agent and then collected trajectories with an expert policy perturbed by uniform noise spanning $[0.3,0.7]$ to ensure adequate coverage of the state-action space. We trained five comparison models; all share the same data scale (128K trajectories $\times$ 200 step/trajectory) but differ in the number and scope of the environments.
@@
Figure 2: Comparison of models trained on different datasets (color-coded) in Cart-Poles. Performance varies markedly with the training data, revealing distinct tendencies toward ER, EL, or an inability to perform ICL.
@@
Importance of both environment scope and environment number: A comparison between the model trained on 1 environment (1 Env) and 4 environments (4 Envs) with the other group demonstrates that an insufficient number of environments leads to the absence of ICL and generalization, except in the tasks that the model has already seen. The 4 Envs group exhibits clear ER characteristics, with a substantial performance gap between seen and unseen tasks. The 16 Envs (Scope 1 + 2) and 8K Envs (Scope 1) groups display similar capabilities; however, they lag significantly behind the 8K Envs (Scope 1 + 2) group, indicating that both the scope of tasks and the number of tasks are crucial.
@@
Figure 3: Best Matching Error (BME) versus prediction error for various models across 130 test cart-poles.
EL exhibits a smaller generalization gap and a relatively lower upper error bound. According to Equation˜6, as the context length increases, the Best Matching Error (BME) remains within the upper error bound in ER, thereby limiting generalization. In contrast, EL is not affected by this term. To further investigate these differences on a case-by-case basis, we examine the correlation between the BME of each of the 130 unseen test cart-poles and their corresponding prediction errors. The BME for each model and testing environment is estimated by applying the ground-truth world model of each environment in the training set to the test environment and selecting the minimum error. Note that we plot the prediction error for $T>100$; therefore, the terms containing $T^{-1/2}$ are negligible, reflecting asymptotic performance. The results, shown in Figure˜3, reveal that the model trained with only four environments exhibits upper error bounds close to the line $\text{error}=\text{BME}$, whereas the upper error bound progressively moves below this line as the number of training environments increases. This not only validates the BME as a crucial term in Equation˜6 but also confirms a transition from ER to EL mode as the number of environments increases.
@@ 4.2 Navigation
Experiment settings. Procedurally generated mazes are a demanding test-bed for transition prediction (Pašukonis et al., 2023; Wang et al., 2024a). By stripping observations of semantic cues, the maze framework keeps task complexity low while still exposing models to stochastic, partially observable dynamics. We amplify diversity through fully randomized configurations that vary topology, textures, object placement, and agent embodiment  [^1] Procedural Maze environments:
@@
. Room-tour data are collected in a way similar to cart-pole: we perturb the oracle object-navigation policy (Ehsani et al., 2024; Pašukonis et al., 2023) with random noise levels in $[0.05,0.95]$ and record the complete trajectory of an agent in each environment. The oracle policy was derived using Dijkstra’s algorithm based on the ground truth 2-D occupancy maps. Observations are RGB images standardized to 128 × 128 pixels; The action space comprises 17 discrete actions, each corresponding to a unique offset and rotational movement. Table 1 lists the resulting training sets, each drawn from a different number of environments and exhibiting varying trajectory lengths. To isolate the effect of data distribution, all maze datasets contain the same number of frames but differ in the coverage of trajectories and environments. To further investigate transferability to more realistic environments, we also collect trajectories from the semantically rich ProcTHOR simulation, which offers a wide variety of assets Kolve et al. (2017); Deitke et al. (2022). Specifically, we curate two datasets: a larger one with 40,000 trajectories and a smaller one with 5,000 trajectories, each trajectory having a length of 2,000 frames.
@@
For training, inspired by overshooting Hafner et al. (2019b) and Hu et al. (2022), we randomly mask the $s_{t}$ of the input sequences at some positions to enhance the model’s capability to predict the distant future (see training details in Section˜B.1). Evaluation is conducted on both seen and unseen tasks. By default, we use an evaluation set scale of $|\mathcal{E}|=256$. The evaluation process involves encoding a context of length $t$ for EL and then predicting future $k$-step transitions using auto-regression and the ground truth action records $a_{t+1},a_{t+2},...,a_{t+k}$. We assess the error in both the latent spaces and the decoded images, which we refer to as $k$-step prediction with context length $T=t$. For example, $T=10$ and $k=4$ predict future $4$ steps with a context length of $10$.
We evaluate two additional baselines alongside the proposed long-context world model: (1) Navigation World Model (NWM) Bar et al. (2025), which employs diffusion layers to predict the next frame from the preceding four frames; (2) Dreamer-v3 Hafner et al. (2019a), which uses LSTM layers for temporal encoding. For NWM, we retain its original pre-trained image encoders and re-train only the diffusion layers on the target dataset. For Dreamer-v3, we remove the policy components and train only the world-model module to ensure a fair comparison.
@@
Impact of data distribution and model architecture on ICL. Table˜2 reports the next-frame prediction quality estimated by PSNR ($k=1$) for models trained on different datasets. Three empirical findings corroborate Theorem˜1: (1) The 32K-L dataset yields the best generalization to unseen environments, whereas the 128-L dataset excels on seen ones; in both cases, peak performance occurs at the asymptotic stage, not at the beginning of the context. (2) Long-context training consistently produces stronger ICL than short-context training, confirming that extensive context is necessary for ICL to emerge. (3) Dreamer and NWM fall short even with a long-context dataset: Dreamer’s LSTM backbone and NWM’s 4-frame horizon show that architectures incapable of fully leveraging long contexts cannot achieve many-shot ER. Figure˜4 further presents the $k=\{1,2,4,8\}$-step prediction performances on unseen Mazes, measuring how far ahead the world models can reliably foresee. For $k>1$, the performance–context-length curve largely tracks the next-frame trend across models, except for Dreamer (Maze 32K-L): at $k=8$, its performance plateaus once $T>100$, whereas $k=1$ keeps improving, revealing a larger compound-error accumulation than in our method.
@@
Figure 5: The decline in performance of EL (trained with Maze-32K-L) and ER (trained with Maze-128-L) when observations in contexts are shuffled, measured by PSNR.
EL transfers better than ER. In Table˜3, we train L2World on ProcTHOR trajectories and evaluate it on unseen ProcTHOR scenarios. The EL model pre-trained on Maze-32K-L not only excels in unseen mazes but also maintains its advantage when fine-tuned on the small ProcTHOR-5K dataset. Its transferability significantly surpasses that of Maze-128-L and other baselines, demonstrating EL’s domain generality. Further increasing the amount of ProcTHOR data leads to continuous improvement in our model while preserving a substantial margin over Dreamer and NWM. However, performance at $T=1K$ to $T=10K$ begins to deteriorate when the ProcTHOR training data increases from 5K to 40K, suggesting that insufficient-length data ($T\leq 2K$) impairs the long-ICL ability acquired in Maze scenarios.
EL is more sensitive to context perturbations than ER. We investigate how models trained on datasets with varying levels of diversity (Maze-32K-L versus Maze-128-L) respond to perturbations in context. Specifically, to assess the importance of context on performance, we randomly shuffle 20% or 50% of the observations within contexts while keeping the actions unchanged. The results are illustrated in Figure˜5. Interestingly, we find that models trained on Maze-32K-L (which are expected to exhibit EL) are more severely affected by these perturbations than those trained on Maze-128-L. This suggests that EL depends more heavily on context, whereas ER relies more on model parameters and is therefore less influenced by changes in context.
@@ 5 Conclusions and Limitations
Conclusions: This work investigates in-context learning of world models, specifically dynamic models, focusing on the possible modes of EL and ER in MDP and POMDP. Theoretically, we analyze error upper bounds for both modes to characterize their properties and identify the conditions under which each excels. Empirically, we introduce L2World and validate these insights in cart-pole and navigation tasks. Our results underscore that both high environment diversity and sufficient context length of the world model are essential to elicit EL.
@@
Acknowledgments
This project is supported by Longgang District Shenzhen’s “Ten Action Plan” for supporting innovation projects (Grant LGKCSDPT2024002), Guangdong Provincial Leading Talent Program (Grant No. 2024TX08Z319), National Natural Science Foundation of China (62033012), Guangdong S&T Program (Grant No.2025B0909040003), and the "Zhiguo" Action of Guangxi Science and Technology Program under Grant No.ZG2503980003.
Ethics Statement
This research adheres to the ethical standards of the machine learning community. The datasets used in this work were synthetic and designed solely for academic research; they do not contain personal, sensitive, or identifiable information.
Reproducibility Statement
We provide details of datasets and experiment settings in Section˜4 and Appendix˜B. The training and evaluating codes are available at https://github.com/airs-cuhk/airsoul/tree/main/projects/MazeWorld. The benchmarking environments are available at https://github.com/FutureAGI/Xenoverse/blob/main/xenoverse/metacontrol/random_cartpole.py (random cart-pole) and https://github.com/FutureAGI/Xenoverse/tree/main/xenoverse/mazeworld (maze).
@@ Appendix A Assumptions and Proofs for Theorem˜1
We first define
@@
Proof for the first part of Theorem˜1: First, we gave that
@@
$=TV(\hat{p}_{\theta,\hat{e}},\hat{p}_{\theta,\hat{e}_{0}})+\min_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}})$ (7)
@@
$<\frac{\alpha(|\mathcal{E}|-1)}{3\sqrt{T}}$ (8)
On the other hand, by definition, the TV of ER satisfies the following upper bound:
$TV(\hat{p}_{ER},p_{e_{0}})$ $\leq\max_{e_{0}\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}}),$
$\leq\max_{e_{1},e_{2}\in\mathcal{E}}TV(p_{e_{1}},p_{e_{2}})+\min_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}}).$ (9)
By synthesizing Equation˜7, Equation˜8, and Equation˜9, the proof is complete.
@@
To estimate $\hat{p}_{EL}(s^{\prime}|s,a)$, we use Equation˜5 to acquire:
@@
Now employ Equation˜10 and Equation˜11 in the calculation of TV:
@@ B.1 Environments and Datasets
Cart-pole: Figure˜6 lists the scopes of the Cart-Pole environment variants, their t-SNE visualization, and details of the training and evaluation data. All datasets share a trajectory length of 200, which is sufficient for ICL in Cart-Pole variants. The training data comprises a total of 25.6M timesteps to avoid interference from data scale on performance, and the evaluation data contains 205K steps in total.
@@
Mazes: Our maze environment is closely related to the settings described in Pašukonis et al. (2023); Wang et al. (2024b), which are generated on a $15\times 15$ grid world. The distribution of the sampled mazes is shown in Figure˜7 and Figure˜8. The primary distinction between our mazes and those in previous work is the enhanced environmental diversity achieved through randomized configurations, which include the following:
@@
Figure 9: Illustration of the embodiment, the observation, and the trajectories in procedurally generated mazes (a) and ProcTHOR (b).